OpenStudy (anonymous):

PLEASE HELP MEDAL AND FAN 2 MULTIPLE CHOICE QUESTIONS ONLY

ganeshie8 (ganeshie8):

Join BE and apply `triangle exterior angle theorem`

ganeshie8 (ganeshie8):

Or if you're formula oriented simply use below : \(\text{external angle } = \dfrac{\text{(far arc) - (near arc)}}{2} = \dfrac{100-30}{2}\)

OpenStudy (anonymous):

but what did u get as an aswer